OpenStudy (anonymous):

What are the asymptotes of: (y+3^2) over (36) − (x+4^2) over (64) =1. please help me.

OpenStudy (anonymous):

i just need help with a few. please

OpenStudy (tkhunny):

\(\dfrac{y+9}{36} - \dfrac{x+16}{64} = 1\) This seems to be what you have written. Is that what you meant to write? Perhaps this? \(\dfrac{(y+3)^{2}}{36} - \dfrac{(x+4)^{2}}{64} = 1\)

OpenStudy (anonymous):

the second! sorry

OpenStudy (tkhunny):

First, we need to be sure you are able to communicate. Do you know how to fix what you have posted so that it will be what you intend?

OpenStudy (anonymous):

no. how do i do that?

OpenStudy (anonymous):

do i click on equation and put it there?

OpenStudy (tkhunny):

1) Remember your Order of Operations. There are reasons why you studied this. Remember it ALL! 2) Use parentheses liberally. x + 2^2 is \(x + 4\) (x + 2)^2 is \((x + 2)^{2}\) What a difference parentheses make!

OpenStudy (anonymous):

okay. pemdas right?

OpenStudy (tkhunny):

You can use the Equation Builder. It is very helpful. You can also just learn that typing inline is different from handwriting or textbook printing. You have to make sure you write what you mean. You can't be casual about it.

OpenStudy (anonymous):

oh okay

OpenStudy (tkhunny):

All-righty, then. Let's see you present the problem statement AND show your best work.

OpenStudy (anonymous):

thats what i need help on though. I dont understand

OpenStudy (anonymous):

did you mean just write the problem out? or solve it? @tkhunny

OpenStudy (tkhunny):

Yes!

OpenStudy (anonymous):

oh well i tried this problem and got it wrong. can we move on to the next?

OpenStudy (tkhunny):

That's a horrible failure. What's wrong with this one? No point in moving on without this one.

OpenStudy (anonymous):

i need help with homework...

OpenStudy (tkhunny):

I thought that's what we were doing until you gave up. Would you like to continue?